OpenStudy (anonymous):

Explain how you would find the exact value of sqrt(2)/sqrt(8)

OpenStudy (anonymous):

\[\frac{ \sqrt{2} }{ \sqrt{8} }\]

OpenStudy (anonymous):

\[\sqrt{8}\] = \[\sqrt{4*2}\] = \[2\sqrt{2}\] so you have: \[\frac{ \sqrt{2} }{ 2\sqrt{2} }\] = \[\frac{ 1 }{ 2 }\]